OpenStudy (pink.1d):

What is the amplitude, period, and phase shift of f(x) = -3 cos(4x + π) + 6?

OpenStudy (michele_laino):

hint: using this identity: \[\Large \cos \left( {4x + \pi } \right) = \cos \left( {4x} \right)\cos \pi - \sin \left( {4x} \right)\sin \pi = - \cos \left( {4x} \right)\] we can rewrite your function as follows: \[\Large f\left( x \right) = 3\cos \left( {4x} \right) + 6\]

OpenStudy (michele_laino):

now, the graph of the function: \[\Large f\left( x \right) = 3\cos \left( {4x} \right) + 6\] is: